OpenStudy (anonymous):

Simplify completely 4x+16/6

OpenStudy (callisto):

Is the expression like this?\[\frac{4x+16}{6}\]

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[\frac{4x+16}{6} = \frac{4(x+4)}{2\times 3}\]Can you simplify it?

OpenStudy (anonymous):

2(x+4)/2 i think

OpenStudy (anonymous):

or over 3

OpenStudy (callisto):

Nope... \[\frac{2\times 2(x+4)}{2\times 3} = \frac{2(x+4)}{3} \]

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

Welcome!